# Elimination Distance to Dominated Clusters

### v1

@@ 4.3 Skeletons for general graphs
In this section, we show that by parameterizing only by $q,k$, and $d$ we can compute the skeleton of unbreakable graphs in time $f(q,k,d)\cdot|G|^{\mathcal{O}(d)}$. This construction follows the work of Bentert et al. [3]. Having access to this algorithm will then be useful for our uniform version of their result.
@@
This lemma resembles and can be proved analogously to [3, Lemma 7].
@@
Assume the existence of a solution $S^{\prime}$ and let $S$ be it skeleton. We guess a set $D$ of at most $d$ vertices in time $O(n^{d})$. (This will play the role of the dominating set for the big component). We then define $S_{1}=G-N[D]$ and $S_{2}=\{v\in N(u)\leavevmode\nobreak\ |\leavevmode\nobreak\ u\in S_{1}\wedge\deg(u)\leq q\}$. If either $|S_{1}|$ or $|S_{2}|$ is bigger than $q$, we terminate this branch. Otherwise, we now claim that $S\subseteq S_{2}$.
To this end, let $C^{\prime}_{0}$ be the large connected component of $G-S^{\prime}$ and assume that the algorithm guessed $D$ as a dominating set for $C^{\prime}_{0}$. Observe that $C^{\prime}_{0}\subseteq N[D]\subseteq N[C^{\prime}_{0}]$. Also, a vertex of $G-N[C^{\prime}_{0}]$ has no neighbor in $C^{\prime}_{0}$, and therefore only has degree at mots $q$. Finally, remember that any vertex $v$ of $S$ must have neighbor $u$ that is not in $N[C^{\prime}_{0}]$ and hence not in $N[G]$. So $u$ is in $S_{1}$ and $u$ has degree at most $q$. So $v$ belongs to $S_{2}$.
@@
Note that Lemma 4.25 applies to skeletons for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. Furthermore, the dominating set $D$ for the large connected component is computed (by brute-force).
@@ 5 Annotated versions for Dominated Cluster Deletion
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Dominated Cluster Deletion problem is to find at most $k$ vertices $S\subseteq G-F$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
The next proposition states that this problem can be solved on specific instances which we call bag graphs.
@@
A graph $G$ is called a $(q,k,d)$-bag graph, if the vertices of $G$ can be partition into two vertex sets $I,E$ called respectively interior vertices and exterior vertices, such that:
  • a connected component of $G[E]$ has at most $12d+$ 1 vertices,
  • a connected component of $G[E]$ has at most $q$ neighbors in $I$, and
  • for every separation $L,R$ of $G$ of order $k$ such that $L\cap R\subseteq I$, we have that either $|L\cap I|\leq q$, or $|R\cap I|\leq q$.
@@
There is an algorithm such that, given a $(q,k,d)$-bag graph $G=(I,E)$, and vertex sets $F,R,B$ of $G$ satisfying $E\subseteq F$ and $|E\cap B|\leq qd$, the algorithm decides whether $G$ is a positive instance to $(k,d)$-Annotated Dominated Cluster Deletion, in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
We prove Proposition 5.3 in Appendix A. Using this, we then prove the following statement via dynamic programming on an unbreakable tree decomposition in Appendix B.
@@
There is an algorithm solving Annotated Dominated Cluster Deletion in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying 2.
Remember that 2 implies that, on the bag graphs that are build during the dynamic programming, some parameter with 1 is bounded, hence we can use Proposition 5.3 these bag graphs. Note that Theorem 5.4 implies Theorem 1.1, by setting $F=\emptyset$ and $R=B=V(G)$. The annotated versions for Elimination Distance to Dominated Clusters is introduced and solved in Appendix C.
@@ 6 Hardness Results
In this section we give the details of the hardness results mentioned in the introduction.
@@
A problem is para-NP-hard with respect to a parameter if there exists an instance of the problem with a fixed parameter such that solving this instance is harder than solving any problem in NP.
@@
A kernel is polynomial if $f$ is polynomial.
[…]
Recall that $c$ denotes the degeneracy of a graph. A graph $G$ is $c$-degenerate if every subgraph $H$ has a vertex of degree at most $c$ (in $H$).
@@ 7 Conclusion
We have studied the Dominated Cluster Deletion problem and resolved the open question by Bentert et al. [3] whether the problem is fixed-parameter tractable with respect to the parameters $k+d+c$, by proving uniform fixed-parameter tractability even for the smaller parameter $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph. We also introduced the more general Elimination Distance to Dominated Clusters problem and almost fully classified its parameterized and kernelization complexity. The most interesting missing part of the classification is whether the case $d=0$, which corresponds to TreeDepth, is NP-complete on graphs of bounded maximum degree. We conjecture that this is the case. Our proofs combine the main tools for addressing cut problems, the decomposition theorem into unbreakable parts by Cygan et al. [11] with the main tools for addressing domination type problems [20].
@@ Appendix A Algorithm on bag graphs
This section is devoted to the proof of Proposition 5.3.
@@
It closely follows from the proof of Theorem 4.15 in Section 4. The main differences are that we now work on an annotated version of the problem, and that the graph might not be unbreakable, only the set of interior vertices are (see Definition 5.2).
@@
Let $G=(I,E)$ be a $(q,k,d)$-bag graph, $F$ a set with $E\subseteq F$, $R$ and $B$ additional colors on bag vertices. We call a set $S$ a skeleton for annotated dominated cluster deletion with parameters $k$ and $d$ if there is a superset $S^{\prime}\supseteq S$ of vertices satisfying:
@@
  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in a component $C$ of $G-S$ where $|C\cap I|\leq q$, and $C$ contains at least one Red vertex.
We now define an operation that will be necessary to obtain a statement analogous to Lemma 4.6.
@@
For an integer $i$, we define the $q$-saturation at distance $i$ $\mathsf{sat}_{q}^{i}(X)$ of a vertex set $X$ as $\mathsf{sat}_{q}(X)$ if $i=1$, and $\mathsf{sat}_{q}(\mathsf{sat}_{q}^{i-1}(X))$. I.e. we apply the $q$-saturation operator $i$ times.
Intuitively from a bag vertex, we take all the bag vertices that are adjacent or connected through component vertices, but only if there are few of them (at most $q$).
@@
Let $q,k,d$ be integers, $G$ be a $(q,k)$-bag graph that is a positive instance to Annotated Dominated Cluster Deletion, then $G$ has a $3qd$ Red-Blue dominating set.
@@
Let $(I,E)$ be the partition of $G$, and $S$ be a solution i.e. every connected component of $G-S$ has a $d$-Red-Blue dominating set. Let $C_{0}$ be the only connected component of $G-S$ that contains more than $q$ vertices of $I$. In $G-C_{0}$, there are at most $q$ interior vertices, which together with the $qd$ Red exterior vertices amount to at most $2qd$ Red vertices in $G-C_{0}$. So the Blue dominating set for $G-C_{0}$ is of size at most $2qd$ as well. Together with the $d$ Blue vertices dominating all Red vertices of $C_{0}$, we have a $3qd$ Red-Blue dominating set.
Lemma A.5 (analogous to Lemma 4.6).
@@
If there is a vertex of $S$ in $X$, we are done as $X\subseteq\mathsf{sat}_{q}^{q}(X)$. Otherwise, let $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, let $v$ be a vertex of $S$ and $C$ the connected component of $G-S^{\prime}$ witnessing that $v\in S$ i.e. $|C\cap I|\leq q$, $C$ contains a Red vertex, and $v$ has a neighbor is $C$.
@@
The colored skeleton-algorithm with parameters $q,k,d$ over a bag graph $G$ returns a family of sets of vertices, each set of size at most $k$.
First, it computes a dominating set $X$ of size at most $3qd$ (returning the empty family if there is no such set). Second, it computes the set $\mathsf{sat}^{q}_{q}(X)$. Third, it calls the algorithm with parameters $q,k-1,d$ on $G$ (and get a family $\mathcal{F}_{0}$), and on $G-v$ for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ (and get a family $\mathcal{F}_{v}$). And finally, for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ add $v$ to every set in $\mathcal{F}_{v}$, and return the union of all the families.
@@
Note that for any interior vertex $v$, $G-v$ is a $(q,k-1,d)$-bag graph. Furthermore, if the input graph is a positive instance of Annotated Dominated Cluster Deletion, and if the algorithm guesses correctly a vertex $v$ of a solution, then $G-v$ is also a positive instance to Annotated Dominated Cluster Deletion with parameter $(k-1,d)$. So the algorithm can repeat until it guesses a skeleton $S$ of size at most $k^{\prime}$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 1.
@@
Note that by definition, there is at most one component $C_{0}$ of $G-S$ that contains more than $q$ interior vertices. So in $G-C_{0}$, there are at most $3qd+q$ many Red vertices ($3qd$ for the exterior vertices and $q$ for the interior ones). What remains to be done is hence the following.
@@
    Each such possibility yields a value $k^{\prime\prime}$ of vertices that need to be deleted. If $k^{\prime\prime}>k-k^{\prime}$ we reject the instance (in this branch). Remember that $k^{\prime}=|S|$.
@@
    Test whether an additional set $W\subseteq C_{0}-F$ of $k^{\prime\prime\prime}=k-k^{\prime}-k^{\prime\prime}$ vertices that can be deleted from $C_{0}$ such that $G[C_{0}-W]$ has a $d$-Red-Blue dominating set. Hence, the remaining problem is an instance of Partial Red-Blue Domination with parameter $k^{\prime\prime\prime},d$.
    If $G$ is a positive instance of Annotated Dominated Cluster Deletion, witnessed by some set $S^{\prime}$, and if $S$ is a skeleton of $S^{\prime}$, such a set $W$ exists (as $S^{\prime}\cap C_{0}$ is a candidate).
    If $t(\cdot)$ is a parameter satisfying 1, we can then test for the existence of such $W$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found, we reject the instance (in this branch).
This concludes the proof of Proposition 5.3 and therefore Appendix A.
Appendix B Algorithm through unbreakable tree decomposition.
In this section, we assume that we have a fix graph $G$, and $(T,\mathsf{bag})$ a $(q,k)$-unbreakable tree decomposition of $G$. See Section 2 for the definition of an unbreakable tree decomposition, and Theorem 2.1 ([12, Theorem 10]) for the proof of its existence.
The goal of this section is to prove Theorem 1.1 using the fact that Annotated Dominated Cluster Deletion is FPT on bag graphs (Proposition 5.3).

### v2

@@ 4.3 Skeletons for general graphs
In this section, we show that by parameterizing only by $q,k$, and $d$, we can compute the skeleton of unbreakable graphs in time $f(q,k,d)\cdot|G|^{\mathcal{O}(d)}$. This construction follows the work of Bentert et al. [4]. Having access to this algorithm will then be useful for our uniform version of their result.
@@
This lemma resembles and can be proved analogously to [4, Lemma 7].
@@
Assume the existence of a solution $S^{\prime}$ and let $S$ be its skeleton. We guess a set $D$ of at most $d$ vertices in time $O(n^{d})$. (This will play the role of the dominating set for the big component). We then define $S_{1}=G-N[D]$ and $S_{2}=\{v\in N(u)\leavevmode\nobreak\ |\leavevmode\nobreak\ u\in S_{1}\wedge\deg(u)\leq q\}$. If either $|S_{1}|>q$ or $|S_{2}|>q+kq$, we conclude that this $D$ is not a good candidate for the dominating set of the large connected component and terminate this branch. Otherwise, we now claim that $S\subseteq S_{2}$.
To this end, let $C^{\prime}_{0}$ be the large connected component of $G-S^{\prime}$ and assume that the algorithm guessed $D$ as a dominating set for $C^{\prime}_{0}$. Observe that $C^{\prime}_{0}\subseteq N[D]\subseteq N[C^{\prime}_{0}]$. A vertex of $G-N[C^{\prime}_{0}]$ has no neighbor in $C^{\prime}_{0}$, and therefore only has degree at most $q$. Finally, remember that any vertex $v$ of $S$ must have a neighbor $u$ that is not in $N[C^{\prime}_{0}]$ and hence not in $N[D]$. So $u$ is in $S_{1}$ and $u$ has degree at most $q$. So $v$ belongs to $S_{2}$.
Regarding the cardinality constraints. Note that at most $q$ vertices are in $G-C^{\prime}_{0}$ and hence at most $q$ are in $G-N[D]=S_{1}$. Also, the only vertices of $S_{2}$ that are in $C_{0}$ must be neighbors of a vertex of $S^{\prime}$ of degree at most $q$, hence $|S_{2}\cap C_{0}|\leq kq$, and since $|G-C_{0}|\leq q$, we get $|S_{2}|\leq q+kq$.
@@
Note that Lemma˜4.25 applies to skeletons for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. Furthermore, the dominating set $D$ for the large connected component is computed (by brute-force).
@@ 5 Annotated versions for Dominated Cluster Deletion
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Dominated Cluster Deletion problem is to find at most $k$ vertices $S\subseteq G-F$ such that every connected component $C$ of GSG-S has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
The next proposition states that this problem can be solved on instances we call bag graphs.
@@
A graph $G$ is called a $(q,k,d)$-bag graph, if the vertices of $G$ can be partitioned into two vertex sets $I,E$ called respectively interior vertices and exterior vertices, such that:
  • every connected component of $G[E]$ has at most $2d+1$ vertices,
  • every connected component of $G[E]$ has at most $q$ neighbors in $I$, and
  • for every separation $L,R$ of $G$ of order $k$ such that $L\cap R\subseteq I$, we have that $|L\cap I|\leq q$ or $|R\cap I|\leq q$.
@@
There is an algorithm such that, given a $(q,k,d)$-bag graph $G=(I,E)$, and vertex sets $F,R,B$ of $G$ satisfying $E\subseteq F$ and $|E\cap B|\leq qd$, the algorithm decides whether $G$ is a positive instance to $(k,d)$-Annotated Dominated Cluster Deletion, in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
We prove Proposition˜5.3 in Appendix˜A. Using this, we then prove the following statement via dynamic programming on an unbreakable tree decomposition in Appendix˜B.
@@
There is an algorithm solving Annotated Dominated Cluster Deletion in time $f(q,k,d,t(G))\cdot|G|^{O(1)}$, where $t(\cdot)$ is a parameter satisfying ˜2.
Remember that ˜2 implies that some parameter with ˜1 is bounded on bag graphs built in our algorithm, so we can use Proposition˜5.3 on these bag graphs. Note that Theorem˜5.4 implies Theorem˜1.1, by setting $F=\emptyset$ and $R=B=V(G)$. The annotated version for Elimination Distance to Dominated Clusters is introduced and solved in Appendix˜C.
@@ 6 Hardness Results
In this section, we give the details of the hardness results mentioned in the introduction.
@@
A problem is para-NP-hard with respect to a parameter if the restriction to instances whose parameter value is at most a certain constant, is itself an NP-hard problem.
@@
A kernel is polynomial if $f$ is. We use the AND-cross-composition technique [19] to show that Elimination Distance to Dominated Clusters does not admit a polynomial kernel unless $\textsf{NP}\subseteq\textsf{coNP/poly}$. If an NP-hard language $L$ AND-cross-composes into a parameterized language $Q$, then $Q$ does not admit a polynomial kernel, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$. Since the treedepth of a disjoint union of a family of graphs is equal to the maximum over the treedepth of these graphs, the disjoint union yields an AND-cross-composition from the unparameterized version of Treedepth into the parameterized one. As computing the treedepth of a graph is NP-hard [32], it follows that Treedepth not admit a polynomial kernel, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$. It remains to prove that Treedepth is also NP-complete on $2$-degenerate graphs. Recall that $c$ denotes the degeneracy of a graph. A graph $G$ is $c$-degenerate if every subgraph $H$ has a vertex of degree at most $c$ (in $H$).
@@ 7 Conclusion
We have studied the Dominated Cluster Deletion problem and resolved the open question by Bentert et al. [4] whether the problem is fixed-parameter tractable with respect to the parameters $k+d+c$, by proving uniform fixed-parameter tractability even for the smaller parameter $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph. We also introduced the more general Elimination Distance to Dominated Clusters problem and almost fully classified its parameterized and kernelization complexity. The most interesting missing part of the classification is whether the case $d=0$, which corresponds to Treedepth, is NP-complete on graphs of bounded maximum degree. We conjecture that this is the case. Our proofs combine the main tools for addressing cut problems, the decomposition theorem into unbreakable parts by Cygan et al. [12] with the main tools for addressing domination-type problems [20].
@@ Appendix A Algorithm on bag graphs
This section is devoted to the proof of Proposition˜5.3.
@@
It closely follows from the proof of Theorem˜4.15 in Section˜4. The main differences are that we now work on an annotated version of the problem, and that the graph might not be unbreakable, only the set of interior vertices are (see Definition˜5.2).
@@
Let $G=(I,E)$ be a $(q,k,d)$-bag graph, $F$ a set with $E\subseteq F$, $R$ and $B$ additional colors on interior vertices. We call a set $S$ a skeleton for annotated dominated cluster deletion with parameters $k$ and $d$ if there is a superset $S^{\prime}\supseteq S$ of vertices satisfying:
@@
  • $S$ contains exactly the vertices of $S^{\prime}$ that have at least one neighbor in a component $C$ of GSG-S where $|C\cap I|\leq q$, and $C$ contains at least one Red vertex.
We now define an operation that will be necessary to obtain a statement analogous to Lemma˜4.6.
@@
For an integer $i$, we define the $q$-saturation at distance $i$ $\mathsf{sat}_{q}^{i}(X)$ of a vertex set $X$ as $\mathsf{sat}_{q}(X)$ if $i=1$, and $\mathsf{sat}_{q}(\mathsf{sat}_{q}^{i-1}(X))$, i.e. we apply the $q$-saturation operator $i$ times.
Intuitively from an interior vertex, we take all the interior vertices that are adjacent or connected through exterior vertices, but only if there are few of them (at most $q$).
@@
Let $q,k,d$ be integers and $G$ be a $(q,k,d)$-bag graph that is a positive instance to Annotated Dominated Cluster Deletion, then $G$ has a $3qd$ Red-Blue dominating set.
@@
Let $(I,E)$ be the partition of the $(q,k,d)$-bag graph $G$, and $S$ be a solution of Annotated Dominated Cluster Deletion, i.e. every connected component of GSG-S has a $d$-Red-Blue dominating set. Let $C_{0}$ be the only connected component of GSG-S that contains more than $q$ vertices of $I$. In $G-C_{0}$, there are at most $q$ interior vertices, which together with the $qd$ Red exterior vertices amount to at most $2qd$ Red vertices in $G-C_{0}$. So the Blue dominating set for $G-C_{0}$ is of size at most $2qd$ as well. Together with the $d$ Blue vertices dominating all Red vertices of $C_{0}$, we have a $3qd$ Red-Blue dominating set.
Lemma A.5 (analogous to Lemma˜4.6).
@@
If there is a vertex of $S$ in $X$, we are done as $X\subseteq\mathsf{sat}_{q}^{q}(X)$. Otherwise, let $S^{\prime}\supseteq S$ witnessing that $S$ is a skeleton, let $v$ be a vertex of $S$ and $C$ the connected component of $G-S^{\prime}$ witnessing that $v\in S$, i.e. $|C\cap I|\leq q$, $C$ contains a Red vertex, and $v$ has a neighbor is $C$.
@@
The annotated skeleton-algorithm with parameters $q,k,d$ over a $(q,k,d)$-bag graph $G$ returns a family of sets of vertices, each set of size at most $k$.
First, it computes a dominating set $X$ of size at most $3qd$ (returning the empty family if there is no such set). Second, it computes the set $\mathsf{sat}^{q}_{q}(X)$. Third, it calls the algorithm with parameters q,k1,dq,k-1,d on $G$ (and get a family $\mathcal{F}_{0}$), and on GvG-v for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ (and get a family $\mathcal{F}_{v}$). And finally, for every $v$ in $\mathsf{sat}^{q}_{q}(X)$ add $v$ to every set in $\mathcal{F}_{v}$, and return the union of all the families.
@@
Note that for any interior vertex $v$, GvG-v is a (q,k1,d)(q,k-1,d)-bag graph. Furthermore, if the input graph is a positive instance of Annotated Dominated Cluster Deletion, and if the algorithm guesses correctly a vertex $v$ of a solution, then GvG-v is also a positive instance to Annotated Dominated Cluster Deletion with parameter (k1,d)(k-1,d). So the algorithm can repeat until it guesses a skeleton $S$ of size at most $k^{\prime}$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜1.
@@
Note that by definition, there is at most one component $C_{0}$ of GSG-S that contains more than $q$ interior vertices. So in $G-C_{0}$, there are at most $3qd+q$ many Red vertices ($3qd$ for the exterior vertices and $q$ for the interior ones). What remains to be done is hence the following.
@@
    Each such possibility yields a value $k^{\prime}$ of vertices that need to be deleted. If $k^{\prime}>k-s$ where $s=|S|$, we conclude that this $S$ is not a skeleton because there is no possible $S^{\prime}$.
@@
    Test whether an additional set $W\subseteq C_{0}-F$ of $k^{\prime\prime}=k-s-k^{\prime}$ vertices that can be deleted from $C_{0}$ such that $G[C_{0}-W]$ has a $d$-Red-Blue dominating set. Hence, the remaining problem is an instance of Partial Red-Blue Domination with parameter $k^{\prime\prime}$ and $d$. If $G$ is a positive instance of Annotated Dominated Cluster Deletion, witnessed by some set $S^{\prime}$, and if $S$ is a skeleton of $S^{\prime}$, such a set $W$ exists (as $S^{\prime}\cap C_{0}$ is a candidate). If $t(\cdot)$ is a parameter satisfying ˜1, we can then test for the existence of such $W$ in time $f(q,k,d,t(G))\cdot|G|^{\mathcal{O}(1)}$. If no solution is found, then this $S$ is not a skeleton.
This concludes the proof of Proposition˜5.3 and therefore Appendix˜A.
Appendix B Algorithm through unbreakable tree decomposition
In this section, we assume that we have a fixed graph $G$, and a $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$ of $G$. See Section˜2 for the definition of an unbreakable tree decomposition, and Theorem˜2.1 ([13, Theorem 10]) for the proof of its existence.
The goal of this section is to prove Theorem˜1.1 using the fact that Annotated Dominated Cluster Deletion is FPT on bag graphs (Proposition˜5.3).